A plane flies 720 mi. against a steady 30-mi/hr headwind and then returns to the same point with the wind. If the entire trip took 10hrs, what is the planes speed in still air?
Answer by Scriptor(36)

Hello,
This question is not as innocent as it looks :-)
First thing to know:
SPEED * TIME = DISTANCE

Call T the time to get to the point at 720 miles.
Call t the time to get back at the starting point.
We know that T + t = 10
Call x the speed of the plane. We need to know x.
Using speed*time = distance we obtain:
(x-30)T + (x+30)t = 1440
Replace T by 10 -t and work it out to obtain:
x + 6t = 174
But: we also know that the distance is 720 (for a one way ticket :-) )
Thus: (x + 30)t = 720
Replace x by 174 - 6t to obtain:
-6t² + 204t - 720 = 0
Solve this for t. Hint: D = 156²
You get 2 solution: t= 4 and t=30
t = 30 is false because the time < 10
We now know that t = 4 and thus T = 6
Since (x + 30)t = 720 => (x+30)*4 = 720
x = 150
Greets,
Scriptor